OpenStudy (anonymous):

solve this system of equations problem, both are togethor as the problem 2x-y=12 x+3y=13

OpenStudy (anonymous):

2x-y=12 -2(x+3y=13) ------------ -7y=-14 y=2 2x-2=12 x=7 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . x=13-3y 2(13-3y) - y = 12 26 - 6y - y = 12 -7y = -15 y = 2 x+3(2) = 13 x = 7

OpenStudy (anonymous):

Solving a system of equations means finding their point of intersection. You can find this the most easily by rewriting both equations in slope-intercept form: y=mx+b and then setting them equal to each other. 2X-Y=12 2X-2X-Y=12-2X -Y=12-2X -Y=12-2X ---- ------- -1 -1 Y=-12+2X or Y=2X-12 <---This the equivalent of your first equation X+3Y=13 X-X+3Y=13-X 3Y=13-X 3Y=13-X --- ------ 3 3 Y=4 1/3-1/3X or Y=-1/3X+4 1/3 <---This is the equivalent of the second So now you set them equal to each other like this. 2X-12=4 1/3-1/3X 2X-12+12=4 1/3+12-1/3X 2X=16 1/3-1/3X 2X=16 1/3-1/3X 2X+1/3X=16 1/3-1/3X+1/3X 2 1/3X=16 1/3 X=7 This gives you the X-value of the point of intersection, and all you have to do now is plug this X-value into either of the original (or new) equations to get the Y-value for the coordinate. Sorry for the long answer!